1.) Calculate ΔE if 120kJ of heat energy is absorbed by the system and 40kJ of work





is done on the surroundings.





2.) Find ΔE when a gas absorbs 30kJ of hear and has 15kJ of work done on it.





3.) If the internal energy of a reaction is 480J and the system absorbs 839J of heat





from it’s surrounds,





a. Find the amount of work.





b. Is the work done by the system or work is done on the system? Explain.







Please solve it like this







1. SOLUTION (what formula you use and how you solve it)







2. SOLUTION (what formula you use and how you solve it)







3.A.SOLUTION(what formula you use and how you solve it)





B. EXPLAIN

1
Expert's answer
2022-03-28T17:13:06-0400

The first low of thermodynamics can be given as follows

dE= Q-W

Calculate ΔE if 120kJ of heat energy is absorbed by the system and 40kJ of work

is done on the surroundings.

"\Delta E= 120-(-40)= 160 kJ"


2 . "\Delta E= 30-(15)= 15 kJ"

3 .

 If the internal energy of a reaction is 480J and the system absorbs 839J of heat

from it’s surrounds

"480= 839-(W),W= 359 kJ"

since work is positive so work is done on the system.